### v1

# On the applicability of Kolmogorov’s theory of probability

Abstract
It is a common view that von Neumann laid the foundations of a “non-commutative probability theory” with his axiomatization of quantum mechanics (QM). As such, it is regarded a generalization of the “classical probability theory” due to Kolmogorov. Outside of quantum physics, however, Kolmogorov’s axioms enjoy universal applicability. This raises the question of whether quantum physics indeed requires such a generalization of our conception of probability or if von Neumann’s axiomatization of QM was contingent on the absence of a general theory of probability in the 1920s.
In this work I argue in favor of the latter position. In particular, I show that for non-relativistic $N$-body quantum systems subject to a time-independent scalar potential, it is possible to construct a mathematically rigorous theory based on Kolmogorov’s axioms and physically natural random variables, which reproduces central predictions of QM. The respective theories are distinct, so that an empirical comparison may be possible. Moreover, the approach can in principle be adapted to other classes of quantum-mechanical models.
Part II of this series will address the projection postulate and the question of measurement in this approach.
Keywords:
Geometric quantum theory  -  Foundations of probability theory  -  Quantum potential
Double slit experiment  -  Uncertainty principle
MSC2020:
81P16  -  81S99  -  60A99  -  81P20
1 Introduction
For almost a century now, the theory of quantum mechanics (QM) has stood firm at the core of our modern understanding of the physical microcosm. Still, as evinced by the multitude of interpretations in existence today [freirejuniorOxfordHandbookHistory2022], the theory continues to mystify even experts in the subject. This state of affairs has given QM the idiosyncratic status of what one could call a well-established theory with contested foundations [freirejuniorQuantumDissidentsRebuilding2015, bacciagaluppiQuantumTheoryCrossroads2009].
So far the following aspect of this problem has only received comparatively little attention [szaboCriticalReflectionsQuantum2001, accardiTopicsQuantumProbability1981, accardiNonKolmogorovProbability2009, fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952, nelsonDerivationSchrodingerEquation1966, redeiQuantumProbabilityTheory2007, streaterClassicalQuantumProbability2000, reddigerMadelungPictureFoundation2017, reddigerProbabilisticFoundationNonrelativistic2022, reddigerMathematicalTheoryMadelung2023]: across the empirical sciences, QM and its descendants are the only scientific theories that are not based on the mathematical theory of probability, as developed by Kolmogorov in 1933 [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933]. Since von Neumann’s axiomatization of QM [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927, vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] predates Kolmogorov’s work, it therefore stands to reason, whether quantum physics indeed requires a different notion of probability or if von Neumann’s axiomatization of QM was contingent on the absence of a general, mathematical theory of probability in the 1920s.
In this work I argue in favor of the latter position. In particular, I show that for non-relativistic $N$-body quantum systems subject to a time-independent scalar potential, it is possible to construct a theory based on Kolmogorov’s axioms and physically natural random variables, which reproduces central predictions of QM (Lem. 1 and Thm. 1). The respective theories are distinct, so that their predictions can in principle be compared empirically.
The main line of reasoning is as follows: First and with a historical focus, I elaborate on the general relationship of mathematical probability theory and QM (Sec. 2). In Sec. 3, a probabilistic view on the double slit experiment and quantum theory as a whole is provided, in line with the ensemble interpretation of QM [ballentineStatisticalInterpretationQuantum1970, pechenkinStatisticalEnsembleInterpretation2022]. This physically motivates the implementation of Kolmogorov’s axioms in quantum theory via the Born rule for position (Sec. 4, in particular Lem. 1). Corresponding observables can be obtained from Madelung’s reformulation of the Schrödinger equation (Sec. 5), in line with [reddigerMadelungPictureFoundation2017]. While this suggests that the theory can be erected upon the Madelung equations in conjunction with the Born rule, a recent review [reddigerMathematicalTheoryMadelung2023] showed that the mathematical theory of the Madelung equations is still underdeveloped. Hence and at least for the time being, another approach is needed. Instead, one may develop a “hybrid theory”, in which the dynamics is borrowed from QM and the basic quantities are rigorously defined in the context of said dynamics. For the aforementioned class of quantum systems this is done in Sec. 6. In result, an alternative theory based on Kolmogorov’s axioms is indeed obtained (Thm. 1).
2 Probability and quantum mechanics
The two modern mathematical theories of probability were developed by the Hungarian mathematician John von Neumann (1903–1957) [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927, vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] and the Russian mathematician Andrej N. Kolmogorov (1903–1987) [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933]. That today there are two such theories, points at a potential discrepancy in the modern foundations of probability theory.
Historically, it is known that both researchers knew each other and followed each other’s work [monastyrskiiJohnNeumann2004]. Still, there does not seem to be any documented evidence of a discussion among the two about this particular topic. [^1] Neither an explicit search nor a consultation of [vonplatoCreatingModernProbability1994] yielded any result. Parthasarathy [parthasarathyObituaryAndreiNikolaevich1988] writes that Kolmogorov showed little interest in QM, adding that “it is a blessing to have enough ground left for future mathematicians to till.”
It would go beyond the scope of this article to provide a direct comparison of the respective theories. Here I only note that in the contemporary literature this conflict is resolved by showing that von Neumann’s “quantum probability theory” may be viewed as a generalization of “classical probability theory,” as discovered by Kolmogorov [accardiTopicsQuantumProbability1981, streaterClassicalQuantumProbability2000, redeiQuantumProbabilityTheory2007, szaboCriticalReflectionsQuantum2001].
While this accords with the general philosophy that QM is to be viewed as a generalization of “classical mechanics”, the particular historical context of the development of the two theories may be put forward as an argument against this conventional point of view.
Mostly guided by the ideas of Bohr and Heisenberg, von Neumann developed his theory of probability with the explicit goal of axiomatizing QM. That – and not the development of a probability theory – was his aim. Indeed, in his most important article on the topic [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927] von Neumann noted that he was merely formalizing “the common calculus of probability.” Clearly, he was not aware that one day his notion of probability would be in conflict with what is generally associated with that term.
Kolmogorov, on the other hand, took a direct approach to the question via measure theory. His fundamental work [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933] was published in 1933. This is one year after von Neumann’s famous book on the mathematical foundations of QM [vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] was published and a total of six years after von Neumann’s article on the topic [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927] appeared.
The common view in the literature therefore implicitly claims, that von Neumann generalized Kolmogov’s theory before the latter could even publish his approach.
One should add to this, that today Kolmogorov’s theory of probability is successfully employed in virtually all of the sciences—except quantum theory. Whereas von Neumann’s axiomatization is employed in quantum theory only. [^2] In [accardiNonKolmogorovProbability2009] supposed counterexamples are provided. Though it may be appropriate to discuss them point-by-point, I shall leave the reader with the remark that I find none of the given examples convincing.
Therefore, the hypotheses I put forward here are that, one, von Neumann’s axiomatization was contingent on the absence of a general, mathematical theory of probability at the time and, two, that the correct theory of probability is indeed the one discovered by Kolmogorov.
If this is the case, then contemporary quantum theory, both non-relativistic and relativistic, requires a revision with regards to the underlying probability theory.
In order to put these hypotheses to scientific scrutiny, the question of whether an empirically successful quantum theory can be constructed on the basis of Kolmogorov’s axioms is to be investigated. In what follows I shall show, that this may indeed be possible.
3 Physical considerations
Before addressing any mathematical details, I shall provide a brief introduction to the central physical ideas.
For this purpose, I begin with a discussion of the double slit experiment with material particles. [^3] I do not claim, that the discussion here applies to electromagnetic radiation as well. See Planck’s remark in [einsteinUeberEntwicklungUnserer1969]. Its character as an “examplar” in the sense of Kuhn [kuhnStructureScientificRevolutions2012] was also recognized by Feynman, who wrote that “it has in it the heart of quantum mechanics” and that “[i]n reality, it contains the only mystery” [feynmanFeynmanLecturesPhysics1963].
For the sake of argument, we shall assume that in the experiment only a single electron is released at a time, so that one can view the runs as mutually independent. Then, if an electron impinges on the detector screen, it creates a single, pointlike impact. Yet, as one repeats the experiment a large number of times, an interference pattern begins to emerge.
In the literature one commonly encounters the claim, that this pattern cannot result from particles in the usual sense of the word [feynmanFeynmanLecturesPhysics1963]. The argument is that the pattern is not the same as the conjunction of the two patterns one obtains with one of the slits closed. That is, if $\sigma_{\text{double}}$ is the ($2$-dimensional, time-independent) probability density of the impact locations on the detector screen for the double slit experiment and $\sigma_{i}$ with $i\in\{1,2\}$ are the probability densities for the experiment with one of the slits closed, then we have
$\sigma_{\text{double}}\neq\frac{1}{2}(\sigma_{1}+\sigma_{2})\,.$ (3.1)
Indeed, if the presence of the other slit did not have an influence on the impact location of an individual electron hitting the screen, then in the above equation equality would hold. In the Copenhagen school this is interpreted as an exhibition of the wave-like nature of electrons, [^4] In textbooks one commonly encounters the description in which a wave function (on the detector) is associated to each slit and the pattern is said to emerge from an “interference term” [feynmanFeynmanLecturesPhysics1963]. However, such accounts are heuristic at best, since in QM $\sigma_{\text{double}}$ is obtained from the time evolution of a single initial wave function [kalbermannSingleDoubleslitScattering2002]. whereas the pointlike impacts are interpreted as an exhibition of their particle-like nature. Wave particle duality states that neither view provides a full picture, they are complementary.
However, in the 1960s Landé expressed a different perspective on the matter: according to him [landeQuantumFactFiction1969], the electrons are pointlike particles and it is only their collective, statistical behavior that is wavelike.
Indeed, the attentive reader may have noticed that the above argument does not rule out that electrons are particles, it only rules out certain notions of locality. For instance, an electron could pass through one of the slits, take a turn into the second slit and then go back through the first one before hitting the screen. Or, following de Broglie [bacciagaluppiQuantumTheoryCrossroads2009], one could imagine that the electron is pointlike, yet accompanied by a wave field in its immediate vicinity, which interacts with the other slit and in result changes the electron trajectory [couderSingleParticleDiffractionInterference2006, pucciWalkingDropletsInteracting2018]. At least qualitatively, both of these behaviors could explain Eq. (3.1). Of course, the argument I put forward here is that physically peculiar behavior can statistically lead to Eq. (3.1), it is not to commit to a particular explanation. In this sense, it seems difficult to evade the conclusion that Landé’s criticism of wave particle duality is justified.
What Landé advocated for is a probabilistic understanding of quantum theory: while quantum theory does not predict the positions of the particles, it does predict the respective probabilities. Following Ballentine [ballentineStatisticalInterpretationQuantum1970, pechenkinStatisticalEnsembleInterpretation2022], quantum theory is concerned with ensembles of similarly prepared systems (of particles and radiation). An ensemble is a theoretical abstraction, consisting of an infinite number of samples. Each sample is a particular quantum system in the aforementioned sense. [^5] In the literature the distinction between the ensemble and the sample is often blurred. In particular, it depends on the context, whether the term ‘quantum system’ refers to a sample, an ensemble, or even a class of ensembles. Here I abide by this customary usage of the term. In line with a frequentist understanding of probability [richardvonmisesProbabilityStatisticsTruth1957] and staying on a heuristic level, the statistics of $n$ samples approach the respective probabilities of the ensemble as $n\to\infty$.
In the above example this means that the relative impact frequencies of the electrons hitting the screen after $n$ runs will approach the probability density $\sigma_{\text{double}}$ as $n\to\infty$, appropriately understood. For $\sigma_{1}$ and $\sigma_{2}$ this is also the case, yet they refer to different physical contexts – i.e. different “similarly prepared systems” – and therefore belong to distinct ensembles.
To summarize my argument: at least on a qualitative level, quantum phenomena like the double slit experiment do not conflict with Kolmogorov’s theory of probability.
Rather, the applicability of Kolmogorov’s axioms to this physical domain is to be tested quantitatively. This can only be achieved indirectly by constructing and testing respective physical theories.
Remark 1
Historically there have been many objections to such probabilistic approaches. [^6] See e.g. Refs. [vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955, kochenProblemHiddenVariables1975, puseyRealityQuantumState2012, dieksNeumannImpossibilityProof2017]. In Chap. 5 of [morettiFundamentalMathematicalStructures2019] Moretti has spelled out many of the underlying mathematical theorems. Addressing all of them would go beyond the scope of this article, so I shall focus on those which I deem most relevant in this context.
First, if one follows the common interpretation of the uncertainty principle, then in quantum theory the notion of a point particle – let alone its trajectory – is not a sensible concept. The common interpretation is, however, not uncontested. Already in the 1930s Popper objected to it on the basis of a probabilistic view on quantum theory [popperZurKritikUngenauigkeitsrelationen1934]: [^7] See also §78 in [popperLogicScientificDiscovery1959]. Landé [landeQuantumFactFiction1969] and Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952] argued along similar lines. the assertion that quantum theory is concerned with probabilities implies that Heisenberg’s inequality cannot be interpreted as a statement on an individual particle. On the contrary, it ought to be understood as a probabilistic statement on the respective (quantum-mechanical) standard deviations (for the ensemble). That is, in the ensemble interpretation Heisenberg’s inequality is not a statement on the “localizability” of individual particles and can therefore not be employed as an argument against the particle model.
Second, the most direct result on the applicability of Kolmogorov’s axioms to quantum phenomena, which I am aware of, is the BCHSH inequality [clauserProposedExperimentTest1969, morettiFundamentalMathematicalStructures2019]. This constitutes a generalization of Bell’s theorem [bellEinsteinPodolskyRosen1964], that overcomes the “unrealistic restriction that for some pair of parameters […] there is perfect correlation” [clauserProposedExperimentTest1969]. If one follows the statement that those inequalities are empirically violated, then the random variable in question may not factorize. Whether the approach described here supports this kind of “non-locality” needs to be investigated in the context of a respective dynamical theory [takabayasiDescriptionPauliMatter1957, janossyHydrodynamicalModelWave1966]. $\diamondsuit$
4 Born’s rule as an approach for basing quantum theory on Kolmogorov’s axioms
The Born rule for position provides a natural approach to the construction of a quantum theory on the basis of Kolmogorov’s axioms. [^8] See [reddigerProbabilisticFoundationRelativistic2024] for how this is to be done in the (general-)relativistic domain. In what follows I shall show how this is to be done for an $N$-body quantum system ($N\in\mathbb{N}$).
Consider a normalized (Schrödinger) wave function $\Psi_{0}\in L^{2}(\mathbb{R}^{3N},\mathbb{C})$ on the configuration space $\mathbb{R}^{3N}$ and denote by $\mathcal{B}^{*}(\mathbb{R}^{3N})$ the corresponding $\sigma$-algebra of Lebesgue sets. Then, by Born’s rule, the probability that the bodies are in a configuration lying in $W\in\mathcal{B}^{*}(\mathbb{R}^{3N})$ is
$\mathbb{P}_{0}\left(W\right)=\int_{W}\left\lvert\Psi_{0}\right\rvert^{2}\negthinspace\left(\vec{r}_{1},\dots,\vec{r}_{N}\right)\,\mathrm{d}^{3N}r\ .$ (4.1)
Indeed, normalization of $\Psi_{0}$ is equivalent to $\mathbb{P}_{0}(\mathbb{R}^{3N})=1$. Furthermore, if $t\mapsto U_{t}$ is any $1$-parameter family of unitary evolution operators determining the time evolution of $\Psi_{0}$ via $\Psi_{t}=U_{t}\Psi_{0}$ for $t\in\mathbb{R}$, we can use $\Psi_{t}$ to define $\mathbb{P}_{t}$ in analogy to Eq. (4.1). Then unitarity of $U_{t}$ implies that $\mathbb{P}_{t}(\mathbb{R}^{3N})=1$ for all $t\in\mathbb{R}$. Accordingly, the following holds.
Lemma 1
The tuple $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N}),\mathbb{P}_{t}\bigr{)}$ is a probability space for every $t\in\mathbb{R}$. $\diamondsuit$
From a mathematical point of view, Lem. 1 is all one requires to base any such $N$-body quantum theory on the basis of Kolmogorov’s axioms. By applying standard probability theory [klenkeProbabilityTheoryComprehensive2013], Lem. 1 determines what a random variable is, how their expectation values are defined, how conditional probabilities are obtained, etc.
To obtain a physical theory based on Kolmogorov’s axioms, however, more is required.
The dynamical evolution $t\mapsto U_{t}$ one can generally carry over from QM. The reason is, that this is less difficult to test empirically than the underlying probability theory: a physically incorrect evolution tends to predict physically incorrect emission spectra, scattering cross sections etc., so that in this respect quantum-mechanical models are comparatively trustworthy.
The actual difficulties concern the observables and the status of the projection postulate. In particular, the ‘observables’ ought to be real random variables on $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N})\bigr{)}$ instead of (unbounded, densely defined) self-adjoint operators on $L^{2}(\mathbb{R}^{3N},\mathbb{C}^{k})$. In what follows I shall motivate and rigorously define such random variables on the basis of physical considerations. The projection postulate and the question of measurement shall be addressed via conditional probabilities in Part II of this series.
5 Madelung’s equations
In 1926 Madelung discovered a reformulation of the $1$-body Schrödinger equation, which is reminiscent Newtonian continuum mechanics [madelungAnschaulicheDeutungGleichung1926, madelungQuantentheorieHydrodynamischerForm1927, nelsonDerivationSchrodingerEquation1966, vonrenesseOptimalTransportView2012, reddigerMadelungPictureFoundation2017]. His reformulation can be generalized to the $N$-body Schrödinger equation, which is what I shall consider here (cf. [janossyHydrodynamicalModelWave1969] and §7.1.2 in [hollandQuantumTheoryMotion1993]).
Assuming that the time-dependent wave function $\Psi$ is smooth, we set
$\rho:=\left\lvert\Psi\right\rvert^{2}$ (5.1)
and
$\vec{v}_{a}:=\frac{\hbar}{m_{a}}\operatorname{Im}(\nabla_{a}\Psi/\Psi)$ (5.2)
for any $a\in\{1,\dots,N\}$ and away from the nodes of $\Psi$.
Then, locally, the Schrödinger equation
$\mathfrak{i}\hbar\frac{\partial}{\partial t}\Psi=\sum_{a=1}^{N}-\frac{\hbar^{2}}{2m_{a}}\Delta_{a}\Psi+V\Psi$ (5.3)
with smooth potential $V$ is equivalent to the following system of PDEs:
$m_{a}\left(\frac{\partial}{\partial t}+\sum_{b=1}^{N}\left(\vec{v}_{b}\cdot\nabla_{b}\right)\right)\vec{v}_{a}\equiv\hphantom{abcd}$
$\hphantom{abcd}-\nabla_{a}V+\sum_{b=1}^{N}\frac{\hbar^{2}}{2m_{b}}\,\nabla_{a}\frac{\Delta_{b}\sqrt{\rho}}{\sqrt{\rho}}\ ,$ (5.4a)
$\frac{\partial\rho}{\partial t}+\sum_{a=1}^{N}\nabla_{a}\cdot\left(\rho\,\vec{v}_{a}\right)=0\ ,$ (5.4b)
$m_{a}\frac{\partial v_{a}^{i}}{\partial x_{b}^{j}}-m_{b}\frac{\partial v_{b}^{j}}{\partial x_{a}^{i}}\equiv 0\ .$ (5.4c)
Those $N$-body Madelung equations, Eqs. (5.4), are dynamical equations for the time-dependent particle position probability density $\rho$ and the so called drift (velocity) field $\vec{v}=(\vec{v}_{1},\dots,\vec{v}_{N})$. In particular, the continuity equation, Eq. (5.4b), assures probability conservation and states that the velocity vector field $\vec{v}$ governs the evolution of particle position probability for the ensemble (cf. Sec. 5.1 of [reddigerMadelungPictureFoundation2017]). Eq. (5.4a) is naturally understood as a force equation for the $a$th body and Eqs. (5.4c) constitute a condition of vanishing ‘quantum vorticity’ [reddigerMathematicalTheoryMadelung2023]. The Madelung equations are also employed in de Broglie-Bohm theory [bohmSuggestedInterpretationQuantum1952, bohmSuggestedInterpretationQuantum1952a, hollandQuantumTheoryMotion1993, bohmUndividedUniverseOntological1993] and stochastic mechanics [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952, nelsonDerivationSchrodingerEquation1966, nelsonQuantumFluctuations1985].
The overt analogy to Newtonian continuum mechanics – more specifically, the theory of Brownian motion, as first observed by Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952] – provides us with a number of physically natural random variables.
In this $N$-body theory they are the position of the $a$th body $\vec{r}_{a}$, its (drift) momentum $\vec{p}_{a}=m_{a}\vec{v}_{a}$, its angular momentum $\vec{l}_{a}=(\vec{r}_{a}-\vec{r}_{0})\times\vec{p}_{a}$ around a given $\vec{r}_{0}\in\mathbb{R}^{3}$, and the total energy, as obtained from Eq. (5.4a): [^9] In addition, it is worth considering the stochastic/osmotic velocity for the $a$th body $\vec{u}_{a}=\frac{\hbar}{2m_{a}}\nabla_{a}\rho/\rho$. See e.g. Refs. [nelsonDerivationSchrodingerEquation1966, delapenaDoesQuantumMechanics1982, henkMachineLearningQuantum2023, reddigerMathematicalTheoryMadelung2023].
$E=\sum_{a=1}^{N}\frac{m_{a}}{2}\,\vec{v}_{a}^{2}+V-\sum_{a=1}^{N}\frac{\hbar^{2}}{2m_{a}}\,\frac{\Delta_{a}\sqrt{\rho}}{\sqrt{\rho}}\,.$ (5.5)
Notable derived quantities are the center of mass, the total momentum, and the total angular momentum (around any $\vec{r}_{0}$).
Following this line of reasoning, it may seem like a ‘miracle’ that the following holds: The (Kolmogorovian) expectation value of any one of the aforementioned random variables is exactly equal to the expectation value of the corresponding quantum-mechanical operator. For the $1$-body theory this was first observed by Takabayasi [takabayasiFormulationQuantumMechanics1952]. My contribution to this debate is the recognition that this constitutes evidence for the validity of Kolmogorov’s theory of probability in the quantum domain (cf. [reddigerMadelungPictureFoundation2017] and Sec. 2.3 in [reddigerProbabilisticFoundationNonrelativistic2022]).
In this approach ‘quantization’ still occurs: if one of the aforementioned random variables is a constant, then the induced probability measure is a Dirac measure. This is the case, for instance, for $E$ from Eq. (5.5), whenever $\Psi$ is a stationary state in the quantum-mechanical sense.
Of course, to obtain a mathematically rigorous theory a restriction to local statements and requiring smoothness is not enough. Furthermore, if one ascribes such physical significance to the Madelung equations, then it is natural to build the theory thereon [reddigerMadelungPictureFoundation2017]. To this end, in [reddigerMathematicalTheoryMadelung2023] the mathematical theory of the Madelung equations was reviewed and it was concluded that this theory is still underdeveloped. While Gasser and Markowich did provide a suitable weak formulation [gasserQuantumHydrodynamicsWigner1997], it is not entirely understood, how the Schrödinger equation can be reobtained (for $H^{1}$-wave functions) [reddigerMathematicalTheoryMadelung2023]. [^10] This would also resolve the question of the “classical limit” in a natural manner, see [gasserQuantumHydrodynamicsWigner1997], Rem. 3.1 in [reddigerMathematicalTheoryMadelung2023], and Sec. 5.1 in [reddigerMadelungPictureFoundation2017].
Fortunately, for the purpose of constructing a mathematically rigorous quantum theory on the basis of Kolmogorov’s axioms this does not pose a direct problem. Taking a pragmatic approach, we can carry over the dynamical evolution $t\mapsto U_{t}$ from QM and then only the corresponding physical random variables need to be rigorously defined. This may be understood as a hybrid theory, for it still borrows concepts from QM—as opposed to erecting the entire theory on Madelung’s equations and the Born rule in conjunction with Kolmogorov’s axioms [reddigerMadelungPictureFoundation2017].
6 A hybrid theory
For the purpose of formulating the hybrid theory, we shall constrain ourselves to (time-independent) Hamiltonians of the form
$\hat{H}=\sum_{a=1}^{N}-\frac{\hbar^{2}}{2m_{a}}\Delta_{a}+V\,,$ (6.1)
with $V$ real-valued and Lebesgue measurable. We further require that $\hat{H}$ is self-adjoint on some dense domain $\operatorname{dom}\hat{H}$ in $L^{2}(\mathbb{R}^{3N},\mathbb{C})$ and that $\operatorname{dom}\hat{H}\subseteq H^{2}(\mathbb{R}^{3N},\mathbb{C})$.
Example 1
If $\hat{H}$ is given by Eq. (6.1) and $V$ is a Rellich class potential, then, by the Kato-Rellich theorem [katoFundamentalPropertiesHamiltonian1951], $\hat{H}$ is self-adjoint on $\operatorname{dom}\hat{H}=H^{2}(\mathbb{R}^{3N},\mathbb{C})$. The most prominent class of examples are Coulomb Hamiltonians, which are of central importance in chemistry.
$\diamondsuit$
The above restriction is made, because ‘observables’ always relate to particular models and it is the ‘observables’ relating to (the weak formulation of) Eq. (5.3) we consider here. Nonetheless, the main result in this context, Thm. 1 below, can in principle be generalized to other classes of models, though ultimately physical considerations dictate what random variables one ought to consider. [^11] The definition of the energy in Eq. (6.3c) below is model-independent and an analogous definition may be used for time-dependent Hamiltonians (see e.g. [schmidKatoTheoremIntegration2014], Sec. V in [penzRegularityEvolutionEquations2018], and [balmasedaSchrodingerEquationTimeDependent2022]).
Among the random variables discussed in Sec. 5, only $\vec{v}_{a}$ and $E$ pose mathematical difficulties. $\vec{r}_{a}$ is well-defined and $\vec{l}_{a}$ is defined, once $\vec{v}_{a}$ is. The following theorem solves this problem and shows that the aforementioned equality of expectation values also holds in this more general case.
Theorem 1
Let $\hat{H}\colon\operatorname{dom}\hat{H}\to L^{2}(\mathbb{R}^{3N},\mathbb{C})$ be given via Eq. (6.1) with real-valued and Lebesgue measurable $V$ such that $\hat{H}$ is self-adjoint on a dense domain $\operatorname{dom}\hat{H}\subseteq H^{2}(\mathbb{R}^{3N},\mathbb{C})$ in $L^{2}(\mathbb{R}^{3N},\mathbb{C})$. Furthermore, let $\Psi_{0}\in\operatorname{dom}\hat{H}$ with $\left\langle\Psi_{0}\text{,}\Psi_{0}\right\rangle=1$. For any $t\in\mathbb{R}$ set
$\Psi_{t}=\exp(-\mathfrak{i}t\hat{H}/\hbar)\Psi_{0}\ .$ (6.3a)
Denote the corresponding probability measure by $\mathbb{P}_{t}$ (cf. Lem. 1 above).
Then for all $t\in\mathbb{R}$ the following holds:
  1. 1)
    $\Psi_{t}\in\operatorname{dom}\hat{H}\subseteq H^{2}(\mathbb{R}^{3N},\mathbb{C})$
  2. 2)
    The following Radon-Nikodým derivatives exist and define $L^{1}$-random variables on $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N}),\mathbb{P}_{t}\bigr{)}$:
    $(\vec{v}_{a})_{t}$ $=\frac{\frac{\hbar}{m_{a}}\operatorname{Im}\left(\Psi_{t}^{*}\nabla_{a}\Psi_{t}\right)\mathrm{d}^{3N}r}{\mathrm{d}\mathbb{P}_{t}}$ (6.3b)
    $E_{t}$ $=\frac{\operatorname{Re}\bigl{(}\Psi^{*}_{t}\hat{H}\Psi_{t}\bigr{)}\mathrm{d}^{3N}r}{\mathrm{d}\mathbb{P}_{t}}$ (6.3c)
  3. 3)
    Their expectation values, $\mathbb{E}_{t}(\vec{v}_{a})$ and $\mathbb{E}_{t}(E)$, satisfy the following identities:
    $m_{a}\,\mathbb{E}_{t}(\vec{v}_{a})$ $=\left\langle\Psi_{t}\text{,}-\mathfrak{i}\hbar\nabla_{a}\Psi_{t}\right\rangle$ (6.3d)
    $\mathbb{E}_{t}(E)$ $=\langle\Psi_{t}\text{,}\hat{H}\Psi_{t}\rangle$ (6.3e)
$\diamondsuit$
The idea of using Radon-Nikodým derivatives to define $\vec{v}_{a}$ and $E$ is due to Gasser and Markowich [gasserQuantumHydrodynamicsWigner1997]. A full proof is given in the appendix.
The reader may have noticed that QM makes weaker assumptions on the regularity of wave functions. However, for $L^{2}$-wave functions the existence of the expectation value of e.g. the quantum-mechanical energy is generally not guaranteed, so that even in QM some assumptions of weak differentiability are physically justifiable (cf. Refs. [grublNondifferentiableBohmianTrajectories2011, berryQuantumFractalsBoxes1996, reddigerMathematicalTheoryMadelung2023, carlenConservativeDiffusions1984]). Still, it may be possible to generalize Thm. 1, so that the required order of differentiability on $\Psi_{0}$ is reduced by $1$ (cf. Appx. A in [reddigerMathematicalTheoryMadelung2023]).
Similarly, the fact that $\mathbb{E}_{t}(\vec{r}_{a})$ or $\mathbb{E}_{t}(\vec{l}_{a})$ need not be defined for a given $\Psi_{t}$ does not pose any conceptual problems: they are defined if and only if the analogous quantum-mechanical expectation values are—the respective quantities are equal.
Example 2
We consider the simplest quantum-mechanical model of the hydrogen atom, i.e. $N=1$ and $V(\vec{r})=\text{e}/(4\pi\varepsilon_{0}\left\lvert\vec{r}\right\rvert)$. Using physics conventions, we denote the electron mass by $\mu$ and each stationary state basis function by $\Psi_{nlm}$; $n$, $l$, and $m$ are appropriately chosen integers.
For any such integers we indeed have $\Psi_{nlm}\in\operatorname{dom}\hat{H}=H^{2}(\mathbb{R}^{3},\mathbb{C})$. Recalling that $\Psi_{nlm}(t,\vec{r})\equiv e^{-\mathfrak{i}E_{n}t/\hbar}\,\Psi_{nlm}(\vec{r})$, the respective random variables from Thm. 1.2 are given via
$\vec{v}_{(t,x,y,z)}=\frac{m\hbar}{\mu}\frac{1}{x^{2}+y^{2}}\begin{pmatrix}-y\\ x\\ 0\end{pmatrix}$ (6.4a)
and
$E(t,\vec{r})=E_{n}\ .$ (6.4b)
The angular momentum around the origin, $\vec{l}=\vec{r}\times(\mu\vec{v})$, is not ‘quantized’, yet we may consider the angular momentum around the $z$-axis instead. Denoting the unit vector in $z$-direction by $\vec{e}_{3}$, the latter is given by
$\left(\vec{r}-z\,\vec{e}_{3}\right)\times(\mu\vec{v})=m\hbar\,\vec{e}_{3}\ .$ (6.4c)
Note that, strictly speaking, the aforementioned random variables are only uniquely defined up to a set of measure zero and they are not defined for any $\vec{r}$ with $\Psi_{nlm}(\vec{r})=0$ (also a set of measure zero). Still, there is uniqueness in the sense that there is a maximal domain on which they are continuous; Eqs. (6.4) are to be understood in this manner. $\diamondsuit$
Finally, I would like to remark that the ‘observables’ $\vec{v}_{a}$ and $E$ from Thm. 1 depend on the particular ensemble one considers (as encoded by $\Psi_{t}$). This is in stark contrast to the momentum and energy operator in QM, which generally only depend on the model, not the particular ensemble.
7 Conclusion
It was shown how to construct a mathematically rigorous, non-relativistic quantum theory for an $N$-body system subject to a time-independent scalar potential, which accords with Kolmogorov’s theory of probability. The two main steps were, first, to employ the Born rule for position to construct the probability space (Lem. 1), and, second, to define physically appropriate random variables in relation to the dynamical theory borrowed from QM (Thm. 1).
This approach of constructing a so called hybrid theory can in principle be applied to other quantum-mechanical models, though one does need additional physical arguments to motivate the respective ‘observables’.
With regards to the foundations of mathematical probability theory, it is noteworthy that the equality of the quantum-mechanical and the physically natural ‘Kolmogorovian’ expectation values may be interpreted as evidence for the validity of Kolmogorov’s theory in the quantum domain. [^12] This was already pointed out in [reddigerMadelungPictureFoundation2017]. See also Chap. 2 in [reddigerProbabilisticFoundationNonrelativistic2022]. Thus, Kolmogorov’s axiomatization of probability may indeed be universally applicable. [^13] The physical need for “quantum probability theory” was already put in question by Szabó [szaboCriticalReflectionsQuantum2001], though on the basis of different arguments and a different approach. To my knowledge, the first person to claim the validity of Kolmogorov’s axioms in the quantum domain was Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952].
Of course, it is ultimately an empirical question, whether this approach provides a physically viable competitor to QM. In this respect, differences in prediction are to be explored.
Higher order moments of observables other than position generally differ between the two theories. While equality is again obtained for the energy of any stationary state, it may be worthwhile to consider the energy of quantum systems with a simple time dependence.
A physical problem is that testing such systems generally goes beyond the above model, because the radiation, for instance, needs to be included in the dynamics. In turn, the above argument may need to be generalized to systems with spin. Madelung-type reformulations of the Pauli equation are indeed known [takabayasiDescriptionPauliMatter1957, janossyHydrodynamicalModelWave1966], so that a consideration of the corresponding random variables may provide clues to an experiment.
Finally, the example of angular momentum in Ex. 2 suggests further conceptual differences, some of which may enable more direct experimental inquiries.
[…]
Acknowledgments I would like to thank Bill Poirier, Mike Scherfner, and Max von Renesse for their indirect support of my research efforts. I also thank Jamal Berakdar for helpful comments and Mario Flory for inviting me to a conference, where I held a similarly titled talk. Furthermore, I am indebted to Anhalt UAS for enabling my participation therein.
[…]
Appendix: Proof of Thm. 1
1) The assertion follows directly from the invariance of $\operatorname{dom}\hat{H}$ under time evolution (see e.g. Prop. A.1 in [reddigerMathematicalTheoryMadelung2023]).
2) The two (signed, vector-valued) measures in the “numerator” are well-defined and finite, since $\Psi_{t}\in H^{2}(\mathbb{R}^{3N},\mathbb{C})$. Now, in order to apply the Radon-Nikodým theorem (cf. Cor. 7.34 and Exercise 7.5.1 in [klenkeProbabilityTheoryComprehensive2013]), we must show that for all $W\in\mathcal{B}^{*}(\mathbb{R}^{3N})$ with $\mathbb{P}_{t}(W)=0$ either measure of $W$ at time $t$ also vanishes. For this purpose, take any representative $\Psi_{t}$ of the corresponding $L^{2}$-equivalence class. Then $\mathbb{P}_{t}(W)=0$ implies that $\left\lvert\Psi_{t}\right\rvert^{2}=0$ almost everywhere on $W$ (with respect to the Lebesgue measure; cf. Thm. 4.8 in [klenkeProbabilityTheoryComprehensive2013]). Thus, $\Psi_{t}=0$ almost everywhere on $W$.
3) This holds by construction. $\blacksquare$
[…]